a) 16x 1U 0ll: The quantity of charge q (in C) that passes through a surface varies with time according to the equation qs6 다 3 t-1, (t in s). The instantaneous current (in A) through the surface at a) 49.5 1 s is: b) 127.5 c) 24 d) 7.5 e) 84 Use the folowing informations to answer 012 and 013 twoit ted has a cn re of 10 m and a resistance of 5 f,conected to a potentai difference of 12 volt between its ends 012: The current density (in A/m2 ) is: a) 0.6 x 10 013: The number of electrons that passes through the cross section in t 1s is: a) 0.15 x 10 b) 0.24 x 10 c) 1.2 x 10 d) 0.48 x 10 e) 2.4 x 10 19 19 19 13 b) 1.5 x 10 c) 0.3 x 10 d) 0.375 x 10 e) 0.75 x 10
Answer #1

12.)

By ohm's law ,

V = I*R

where, V = potential difference = 12 volt

R = resistance of rod = 5 ohm

So, current in rod(I) = V/R

I = 12/5

I = 2.4 Amp

Since, current density(J) = I/A

where, A = cross sectional area of rod = 1*10^-6 m^2

So,

J = 2.4/(1*10^-6)

J = 2.4*10^6 amp/m^2

therefore correct option is (e.)

13.)

Since, dQ/dt = I(current)

So,

dQ = I*dt

By integrate,

\int_{0}^{1}1*dQ = \int_{0}^{1}2.4*dt

Q = 2.4*1 - 2.4*0

Q = 2.4 C = passed charge in 1 sec.

Since Q = n*e

So, number of electron passes(n) = Q/e = 2.4/(1.60*10^-19)

n = 1.5*10^19

therefore correct option is (b.)